OpenStudy (anonymous):

if i were to simplify the equation 8x^2 + 2x = 3 , would i divided both sides by 8??

jimthompson5910 (jim_thompson5910):

I would use the quadratic formula

OpenStudy (anonymous):

i have to factor

jimthompson5910 (jim_thompson5910):

8x^2 + 2x = 3 8x^2 + 2x - 3 = 0 8x^2 + 6x - 4x - 3 = 0 Now factor by grouping

OpenStudy (anonymous):

where did you get the 6 and -4 from?

jimthompson5910 (jim_thompson5910):

Multiply 8 and -3 to get -24 Now find two numbers that a) multiply to -24 AND b) add to 2 These two numbers are 6 and -4 So we break up 2x into 6x-4x